# Orthogonal projections of hypercubes

### v3

@@ III.1 PCA by minimizing inner-product error
where $p\left(\bm{s}\right)\in\left[0,1\right]$ is the normalized weight, i.e., the probability of finding the Ising coordinate $\bm{s}$, satisfying $\sum_{\bm{s}}p\left(\bm{s}\right)=1$.
Minimizing $\varepsilon^{2}\right>$ with the normalization constraint $\left\{\bm{v}_{i}^{\top}\bm{v}_{i}=1\right\}_{i=1}^{2}$ corresponds to classical multidimensional scaling (MDS) [119, 120]. Classical MDS666 Also called classical scaling, Torgerson scaling, Torgerson–Gower scaling, or principal coordinates analysis. with squared Euclidean (Pythagorean) distance provides lower-dimensional scaled coordinates preserving the original distance between the vertices. It is shown that classical MDS with squared Euclidean distance is equivalent to PCA [121], which is a linear dimensionality reduction method. See relevant studies [122, 123, 124] for the equivalence of classical MDS and PCA. Interested readers can obtain the derivation of PCA from the minimum inner-product error formulation and other formulations (maximum projection variance formulation and minimum reconstruction error formulation) in Appendix B.
@@
As we show in the following Secs. V.1, VI.2 and VI.3, the projection of hypercubes using PCA is informative because the properties of the projection can be analytically obtained. In addition to the biplot-assisted interpretation of PCA in the next Sec. III.2, we can analytically obtain the properties of the projection of hypercubes for special cases. We analytically reveal both the strengths and weaknesses of PCA in projecting hypercubes in Secs. V.1, VI.2 and VI.3. With such prior knowledge of the projection of hypercubes, we can examine the resulting plots more deeply, even if the knowledge is from special cases.
@@ III.2 PCA and biplot
Considering that PCA is an unsupervised and linear method for projection, we suggest interpreting PCA on binary vertices as a reproducible, interpretable, and automatic projection of a hypercube. PCA [101, 102], a method of statistics and unsupervised machine learning, is an essential technique for analyzing high-dimensional data [103, 104, 125, 126]. By calculating eigenvalues (explained variance) and eigenvectors (PC loadings) of the covariance matrix, PCA finds a set of vectors that maximize the variance of the data777 See Appendix B.1 for the derivation of PCA by the maximum projection variance formulation. and is often employed to perform linear dimensionality reduction.888 In many practical applications of PCA, the probability distribution $p\left(\bm{s}\right)$ is approximated by the empirical distribution of the data points. To perform PCA, one needs to calculate the covariance matrix of the data points. Specifically, in the context of hypercubic vertices, the covariance matrix $\bm{\varSigma}\in\mathbb{R}^{N\times N}$ is defined as
@@
$\bm{e}_{i}=2\begin{bmatrix}u_{i;j}\\ u_{i;k}\end{bmatrix}.$ (10)
From now on, we call $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ biplot basis. Notice that, when the data points are binary vertices $\bm{b}\in\left\{1,0\right\}^{N}$ of a system, the biplot basis exactly match the basis of the orthogonal projection of a hypercube, while when the data points are Ising vertices $\bm{s}\in\left\{+1,-1\right\}^{N}$, the biplot basis with doubled magnitudes exactly match the contribution basis. This difference arises from the difference in the length of hypercubic edges: they are one for binary vertices and two for Ising vertices. This is why we add factor $2$ in Eq. (10). By performing PCA and plotting biplot basis, one can obtain a reproducible, interpretable, and automatic projection of a hypercube.
@@ IV Hypercubic PCA
In the previous Sec. III, we introduced PCA as a method to project hypercubes and presented the interpretation of the projection using biplot basis. PCA, in practice, requires the probability distribution $p\left(\bm{s}\right)$ of the vertices $\bm{s}$ to calculate the covariance matrix. How does $p\left(\bm{s}\right)$ affect the projection of hypercubes? How do the resulting coordinates of the vertices change when $p\left(\bm{s}\right)$ is varied? In this section, we project hypercubes using PCA while varying the probability distribution $p\left(\bm{s}\right)$ of the vertices. Throughout the examples, we reveal the trends of the resulting projections: the leading PC corresponds to the vertices with higher probability, and PCA distorts the distance between the vertices around the origin of the projected space.
Figure 4: Orthogonal projections of four-dimensional hypercubic vertices using PCA. (a) A projection of a four-dimensional hypercube where vertices are weighted randomly. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertex. Arrows are biplot basis and the boxes on the bottom right indicate the correspondence between the colors of the arrows and the original dimensions. (b) Fraction of explained variance by each PC of (a). (c) PC1 loading, and (d) PC2 loading of random weighted hypercubic vertices of (a). (e) Hamiltonian path on a four-dimensional hypercube in (a). For a different realization of random weight, see the Supplemental Material [109]. (f–j) Same as (a–e) but with bipolar distribution. Two of the vertices, $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, are more weighted than the others.
@@ IV.1 Random distribution
The biplot basis in Fig. 4(a) are drawn as arrows from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, assisting us in estimating the original binary coordinates of the vertices. Here, we abbreviate $+1$ to $+$ and $-1$ to --. For example, the projected vertex around $\begin{bmatrix}-2.0&-0.25\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}-&-&+&-\end{bmatrix}^{\top}$ and the projected vertex around $\begin{bmatrix}-0.4&1.7\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$. Equivalently, one can follow the hypercubic edge from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ (the origin of the biplot basis) to the vertex of interest by changing the corresponding elements indicated by the biplot basis to know the original Ising coordinate. Notice that it is not necessary to start from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ to infer the original Ising coordinate: if one knows the original coordinates of any vertex, one can infer the rest from the biplot basis, cf. Eq. (2).
@@
We then present the usage of the projection of the hypercube using PCA. In Fig. 4(e), the Hamiltonian path on the hypercube is shown following the biplot. Starting from the vertex $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, one can follow the path to the vertex $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ in the original high-dimensional space, by knowing which digit changed by following the biplot basis. These properties may be useful for understanding pathways of state transition of the target system. Notice, however, that the arrows indicating the Hamiltonian path overlap around the origin of the projected space. This overlap makes it difficult to follow the path. We address this issue in Sec. VI.
@@ IV.3 Sexapolar distribution
Figure 5: Orthogonal projections of four-dimensional hypercubic vertices using PCA. $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ are the most weighted vertices, $\begin{bmatrix}-&-&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$ are the second most weighted vertices, $\begin{bmatrix}-&+&-&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&-\end{bmatrix}^{\top}$ are the third most weighted vertices, and the rest of them are weighted randomly. (a) Projection of the hypercube by PC1 and PC2. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertices. Arrows are biplot basis and the boxes on the bottom right show the correspondence between the colors of the arrows and the original dimensions. Dashed arrows correspond to the projection of original high-dimensional vectors, $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. The original coordinate is indicated around the lower right of the arrowhead as an array of filled $\blacksquare$ (indicates $+$) or empty $\square$ (indicates --) boxes. Horizontal and vertical dashed lines crossing the origin are for visual aid. (b) Same as (a) but by PC1 and PC3. (c) Fraction of explained variance by each PC. (d) PC1 loading, (e) PC2 loading, and (f) PC3 loading.
@@ IV.4 Brief review and preview
In all orthogonal projections [Figs. 4(a), 4(f), 5(a), and 5(b)] of hypercubes using PCA, the weighted vertices have high PC1 scores, indicating that PC1 is associated with the weighted vertices. This observation is discussed further in Sec. V.
@@ V.1 Analytical investigation
We consider an ideal situation where the distribution is concentrated on a few hypercubic vertices. Consider, for example, a low-temperature canonical ensemble for an Ising spin system with $N$ spins, where the vertex distribution is dominated by a weighted vertex (or ground state) $\bm{\xi}\in\left\{+1,-1\right\}^{N}$ and its globally spin-flipped vertex (state) $-\bm{\xi}$. Suppose the distribution is idealized as a bipolar one
$p\left(\bm{s}\right)\approx\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right),$ (11)
where $\delta_{\bm{y},\bm{x}}$ is the Kronecker delta function for vectors, which is equal to $1$ when $\bm{x}=\bm{y}$ and 00 otherwise. Given that the mean vector is the zero vector, $\bm{s}\right>=\sum_{\bm{s}}\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)\bm{s}=\bm{0},$ the covariance matrix [Eq. (7)] becomes999 If the distribution is unipolar, $p\left(\bm{s}\right)\approx\delta_{+\bm{\xi},\bm{s}},$ the covariance matrix becomes $\bm{\varSigma}=\sum_{\bm{s}}\delta_{+\bm{\xi},\bm{s}}\left(\bm{s}-\bm{\xi}\right)\left(\bm{s}-\bm{\xi}\right)^{\top}=\bm{0}\bm{0}^{\top},$ which results in all vertices being projected to the origin.
$\bm{\varSigma}=\sum_{\bm{s}}\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)\bm{s}\bm{s}^{\top}=\bm{\xi}\bm{\xi}^{\top},$ (12)
which is already diagonalized with the eigenvector $\bm{\xi}$ and the corresponding eigenvalue $N$, i.e., $\bm{\varSigma}\bm{\xi}=\bm{\xi}\bm{\xi}^{\top}\bm{\xi}=N\bm{\xi}$, where $N$ is the largest eigenvalue. Therefore, the PC1 loading is
@@
which satisfies $\bm{\varSigma}\bm{u}_{1}=N\bm{u}_{1}$. Notice the normalization factor $\frac{1}{\sqrt{N}}=\frac{1}{\sqrt{\bm{\xi}^{\top}\bm{\xi}}}$. Also, the element-wise sign of the PC1 loading is the weighted vertex vector, $\operatorname{sgn}\left(\bm{u}_{1}\right)=\bm{\xi}$. This explains why we observe that the element-wise sign of the PC loading is the same as the weighted state. If we remove the normalization factor, the PC loading is exactly same as the weighted state.
The projected coordinates on PC1 (PC1 score) become

### v4

@@ III.1 PCA by minimizing inner-product error
where $p\left(\bm{s}\right)\in\left[0,1\right]$ is the normalized weight, i.e., the probability666 We use the notation $p\left(\bm{s}\right)$ as the probability mass function (discrete probability distribution). of finding the Ising coordinate $\bm{s}$, satisfying $\sum_{\bm{s}}p\left(\bm{s}\right)=1$.
Minimizing $\varepsilon^{2}\right>$ with the normalization constraint $\left\{\bm{v}_{i}^{\top}\bm{v}_{i}=1\right\}_{i=1}^{2}$ corresponds to classical multidimensional scaling (MDS) [108, 109]. Classical MDS777 Also called classical scaling, Torgerson scaling, Torgerson–Gower scaling, or principal coordinates analysis. with squared Euclidean (Pythagorean) distance provides lower-dimensional scaled coordinates preserving the original distance between the vertices. It is shown that classical MDS with squared Euclidean distance is equivalent to PCA [110], which is a linear dimensionality reduction method. See relevant studies [111, 112, 113] for the equivalence of classical MDS and PCA. Interested readers can obtain the derivation of PCA from the minimum inner-product error formulation and other formulations (maximum projection variance formulation and minimum reconstruction error formulation) in Appendix B.
@@
As we show in the following Secs. V.1VI.2 and VI.3, the projection of hypercubes using PCA is informative because the properties of the projection can be analytically obtained. In addition to the biplot-assisted interpretation of PCA in the next Sec. III.2, we can analytically obtain the properties of the projection of hypercubes for special cases. We analytically reveal both the strengths and weaknesses of PCA in projecting hypercubes in Secs. V.1VI.2 and VI.3. With such prior knowledge of the projection of hypercubes, we can examine the resulting plots more deeply, even if the knowledge is from special cases.
@@ III.2 PCA and biplot
Considering that PCA is an unsupervised and linear method for projection, we suggest interpreting PCA on binary vertices as a reproducible, interpretable, and automatic projection of a hypercube. PCA [91, 92], a method of statistics and unsupervised machine learning, is an essential technique for analyzing high-dimensional data [93, 94, 114, 115]. By calculating eigenvalues (explained variance) and eigenvectors (PC loadings) of the covariance matrix, PCA finds a set of vectors that maximize the variance of the data888 See Appendix B.1 for the derivation of PCA by the maximum projection variance formulation. and is often employed to perform linear dimensionality reduction.999 In practical applications of PCA, the probability distribution $p\left(\bm{s}\right)$ is approximated by the empirical distribution of the data points. To perform PCA, one needs to calculate the covariance matrix of the data points. Specifically, in the context of hypercubic vertices, the covariance matrix $\bm{\varSigma}\in\mathbb{R}^{N\times N}$ is defined as
[…]
@@
$\tilde{\bm{e}}_{i}=2\begin{bmatrix}u_{i;j}\\ u_{i;k}\end{bmatrix}.$ (12)
For PCA, we call contribution vector $\tilde{\bm{e}}_{i}$ biplot vector. Notice that, when the data points are binary vertices $\bm{b}\in\left\{1,0\right\}^{N}$ of a system, the biplot basis exactly match the basis of the orthogonal projection of a hypercube, while when the data points are Ising vertices $\bm{s}\in\left\{+1,-1\right\}^{N}$, the biplot basis with doubled magnitudes exactly match the contribution basis. This difference arises from the difference in the length of hypercubic edges: they are one for binary vertices and two for Ising vertices. This is why we add factor $2$ in Eq. (12). By performing PCA and plotting biplot vectors or biplot basis, one can obtain a reproducible, interpretable, and automatic projection of a hypercube.
@@ IV Hypercubic PCA
In the previous Sec. III, we introduced PCA as a method to project hypercubes and presented the interpretation of the projection using biplot vectors. PCA, in practice, requires the probability distribution $p\left(\bm{s}\right)$ of the vertices $\bm{s}$ to calculate the covariance matrix. How does $p\left(\bm{s}\right)$ affect the projection of hypercubes? How do the resulting coordinates of the vertices change when $p\left(\bm{s}\right)$ is varied? In this section, we project hypercubes using PCA while varying the probability distribution $p\left(\bm{s}\right)$ of the vertices. Throughout the examples, we reveal the trends of the resulting projections: the leading PC corresponds to the vertices with higher probability, and PCA distorts the distance between the vertices around the origin of the projected space.
Figure 4: Orthogonal projections of four-dimensional hypercubic vertices using PCA. (a) A projection of a four-dimensional hypercube where vertices are weighted randomly by uniform random numbers in $\left[0,z\right)$. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertex. Arrows are biplot vectors originating from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and the boxes on the bottom right indicate the correspondence between the colors of the arrows and the original dimensions. (b) Fraction of explained variance by each PC of (a). (c) PC1 loading, and (d) PC2 loading of random weighted hypercubic vertices of (a). (e) Hamiltonian path on a four-dimensional hypercube in (a). For a different realization of random weight, see the Supplemental Material [16]. (f)–(j) Same as (a)–(e) but with bipolar distribution. Two of the vertices, $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, are two-times more weighted than the others.
@@ IV.1 Random distribution
The biplot vectors in Fig. 4(a) are drawn as arrows from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, assisting us in estimating the original binary coordinates of the vertices. Here, we abbreviate $+1$ to $+$ and $-1$ to --. For example, the projected vertex around $\begin{bmatrix}-2.0&-0.25\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}-&-&+&-\end{bmatrix}^{\top}$ and the projected vertex around $\begin{bmatrix}-0.4&1.7\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$. Equivalently, one can follow the hypercubic edge from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ (the origin of the biplot vectors) to the vertex of interest by changing the corresponding elements indicated by the biplot vectors to know the original Ising coordinate. Notice that it is not necessary to start from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ to infer the original Ising coordinate: if one knows the original coordinates of any vertex, one can infer the rest from the biplot vectors, cf. Eq. (3).
@@
We then present the usage of the projection of the hypercube using PCA. In Fig. 4(e), the Hamiltonian path on the hypercube is shown following the biplot. Starting from the vertex $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, one can follow the path to the vertex $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ in the original high-dimensional space, by knowing which digit changed by following the biplot vector. These properties may be useful for understanding pathways of state transition of the target system. Notice, however, that the arrows indicating the Hamiltonian path overlap around the origin of the projected space. This overlap makes it difficult to follow the path. We address this issue in Sec. VI.
@@ IV.2 Bipolar distribution
Figure 5: Orthogonal projections of four-dimensional hypercubic vertices using PCA. $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ are the most weighted vertices with weight $3z$, $\begin{bmatrix}-&-&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$ are the second most weighted vertices with weight $2z$, $\begin{bmatrix}-&+&-&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&-\end{bmatrix}^{\top}$ are the third most weighted vertices with weight $z$, and the rest of them are weighted randomly by uniform random numbers in $\left[0,z\right)$. (a) Projection of the hypercube by PC1 and PC2. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertices. Arrows are biplot vectors originating from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and the boxes on the bottom right indicate the correspondence between the colors of the arrows and the original dimensions. Dashed arrows correspond to the projection of original high-dimensional vectors, $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. The original coordinate is indicated around the lower right of the arrowhead as an array of filled $\blacksquare$ (indicates $+$) or empty $\square$ (indicates --) boxes. Horizontal and vertical dashed lines crossing the origin are for visual aid. (b) Same as (a) but by PC1 and PC3. (c) Fraction of explained variance by each PC. (d) PC1 loading, (e) PC2 loading, and (f) PC3 loading.
@@ IV.4 Brief review and preview
In all orthogonal projections [Figs. 4(a), 4(f), 5(a), and 5(b)] of hypercubes using PCA, the weighted vertices have high PC1 scores, indicating that PC1 is associated with the weighted vertices. This observation is discussed further in Sec. V. In the Supplemental Material [16], we show the projections of hypercubes with different realizations of random weight, finding that the vertices does not necessarily have significantly larger weight to have larger PC scores. This result seems to be counterintuitive, but we address this in Sec. V and Appendix C.
@@ V.1 Analytical investigation
We begin with an ideal situation where the distribution is concentrated on a few hypercubic vertices. Consider, for example, a low-temperature canonical ensemble for an Ising spin system with $N$ spins, where the vertex distribution is dominated by a weighted vertex (or ground state) $\bm{\xi}\in\left\{+1,-1\right\}^{N}$ and its globally spin-flipped vertex (state) $-\bm{\xi}$. Suppose the distribution is idealized as a bipolar one as in Fig. 4(f),
$p\left(\bm{s}\right)\approx\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right).$ (13)
Here, $\delta_{\bm{y},\bm{x}}$ is the Kronecker delta function for vectors, which is equal to $1$ when $\bm{x}=\bm{y}$ and 00 otherwise. Given that the mean vector is the zero vector,
$\bm{s}\right>=\sum_{\bm{s}}\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)\bm{s}=\bm{0},$ (14)
the covariance matrix [Eq. (9)] becomes
$\bm{\varSigma}=\sum_{\bm{s}}\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)\left(\bm{s}-\bm{0}\right)\left(\bm{s}-\bm{0}\right)^{\top}=\bm{\xi}\bm{\xi}^{\top},$ (15)
which is already diagonalized with the eigenvector
@@
and the corresponding eigenvalue
$\lambda_{1}=N,$ (17)
i.e., $\frac{1}{\sqrt{N}}\bm{\varSigma}\bm{\xi}=\frac{1}{\sqrt{N}}\bm{\xi}\bm{\xi}^{\top}\bm{\xi}=\frac{1}{\sqrt{N}}N\bm{\xi}$, Notice the normalization factor $\frac{1}{\sqrt{N}}=\frac{1}{\sqrt{\bm{\xi}^{\top}\bm{\xi}}}$. Also, the element-wise sign of the PC1 loading is the weighted vertex vector, $\operatorname{sgn}\left(\bm{u}_{1}\right)=\bm{\xi}$. This explains why we observe that the element-wise sign of the PC loading is the same as the weighted state. If we remove the normalization factor, the PC loading is exactly same as the weighted state.
Next, we consider a finite-temperature canonical ensemble with ground states. Here, the ground states $+\bm{\xi}$ and $-\bm{\xi}$ are weighted more and the other probability of states are approximated as uniform. We assume bipolar distribution with uniform background,
$p\left(\bm{s}\right)\approx c\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)+\left(1-c\right)\frac{1}{2^{N}},$ (18)
where $c\in\left[0,1\right]$ is a parameter that controls the distribution. In the limit $c\to 0$, corresponding to the high-temperature regime, the distribution approaches the uniform distribution, $p\left(\bm{s}\right)\to\frac{1}{2^{N}}$. All states are found randomly in this regime. In the opposite limit, $c\to 1$, the distribution reduces to the bipolar form $p\left(\bm{s}\right)\to\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)$ [Eq. (13)]. Between these two extremes, the distribution can resemble a finite-temperature canonical ensemble, where the ground states are more heavily weighted than the others, but other states also have nonzero probability. For such a distribution, the mean vector remains the zero vector as in the bipolar case, and the covariance matrix becomes
$\bm{\varSigma}=c\bm{\xi}\bm{\xi}^{\top}+\left(1-c\right)\frac{1}{2^{N}}\bm{I},$ (19)
which has the eigenvector $\bm{u}_{1}=\frac{1}{\sqrt{N}}\bm{\xi}$ and the eigenvalue $\lambda_{1}=Nc+1-c$. Thus, as long as $c\neq 0$, the PC1 loading corresponds to that of the bipolar distribution but with a reduced eigenvalue. If $c=0$, the covariance matrix becomes the identity matrix, $\bm{\varSigma}=\bm{I}$, and the eigenvectors can be any set of orthonormal vectors. Therefore, we expect that the PC1 loading remains close to the weighted state as long as the probability distribution retains a bipolar-like structure, i.e., $c>0$. We further discuss this distribution in the next Sec. V.2.
With bipolar or bipolar distribution with background, we move on to the projected coordinates. The projected coordinates on PC1 (PC1 score) become
@@
$r_{1}\left(\bm{s}\right)=\sqrt{N}m\left(\bm{s}\right),$ (21)
as numerically pointed out in previous studies [100, 101, 102].